OpenStudy (umulas):

Can anyone help me with finding cos 3x?

OpenStudy (umulas):

I understand that cos(A+B) = cosAcosB - sinAsinB So I got cos3 cos(2x+1) cos2xcosx-sin2xsinx (2cos^2-1)cosx -2sincosx? How do people get to the part of switching the sinx to cos???? Please help, thanks.!

OpenStudy (amistre64):

cos(x+x) = cos^2 (x) sin^2(x) sin(x+x) = 2 sin(x) cos(x)

OpenStudy (amistre64):

cos(2x) cos(x) - sin(2x) sin(x) (cos^2(x) - sin^2(x)) cos(x) - 2 sin(x) cos(x) sin(x)

OpenStudy (umulas):

But how did you go from the sin part of -sin2xsin into -2sincos I understand that sin2x=2sinxcos, but what happened to the sin in -sin2xsin should it be 2sinxcox(sinx)?

OpenStudy (amistre64):

what is the identity for: sin(x+x)

OpenStudy (amistre64):

oh, you wrote it ...

OpenStudy (amistre64):

since the last part there is\[-sin(2t)~sin(t)\] and sin(2t) = 2sin(t)cos(t), replace it \[-[sin(2t)]~sin(t)\] \[-[2~sin(t)~cos(t)]~sin(t)\] \[-2~sin(t)~cos(t)~sin(t)\]

OpenStudy (amistre64):

forgot it was an x, not a t

OpenStudy (umulas):

(2cos^2-1)cosx -2sin^2cosx 2cosx^3-cosx-2sinx^2cosx. Is this correcg?

OpenStudy (amistre64):

hmmm, id have to collect terms and whatnot (cos^2(x) - sin^2(x)) cos(x) - 2 sin(x) cos(x) sin(x) cos^3(x) - sin^2(x)cos(x) - 2 sin^2(x) cos(x) cos^3(x) - (1-cos^2(x))cos(x) - 2 sin^2(x) cos(x) cos^3 - cos+ cos^3 - 2 sin^2 cos 2cos^3 - cos (1 - 2 sin^2) there are prolly a few different ways to restructure it

OpenStudy (amistre64):

sin^2 = 1-cos^2 2cos^3 - cos (1 - 2 (1-cos^2)) 2cos^3 - cos (1 - 2 + 2 cos^2) 2cos^3 - cos ( - 1 + 2 cos^2) 2cos^3 + cos + 2 cos^3 4cos^3 + cos

OpenStudy (amistre64):

prolly got some errors in there

OpenStudy (amistre64):

\[cos(3 x) = cos^3(x)-3 sin^2(x) cos(x)\]

OpenStudy (amistre64):

in the end after i fix my few erros :) \[cos(3 x) = 4 cos^3(x)-3 cos(x)\]

OpenStudy (umulas):

wait, i have to do more. 2cos^3-cosx-2(1-cosx^2)cosx 2cos^3-cosx-2cosx+2cos^ 4cos^3-3cosx

OpenStudy (umulas):

Thanks for helping.

OpenStudy (amistre64):

youre welcome

OpenStudy (umulas):

Wait a minute.... My answer choices are: cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x So I was wrong?

OpenStudy (anonymous):

we can alternatively come from sinx to cosx by using first principle..do you know it ?

OpenStudy (umulas):

I'm terribly sorry, but I'm confused, may you explain more. By the way, the answer I received was cos x - 4 cos x sin2x ; i just want to know how they got it like that.

OpenStudy (amistre64):

like a stated before, there are a number of different ways to approach this thing, and there is no one standard rewriting of it.

OpenStudy (amistre64):

cos x - 4 cos x sin^2 x cos x - 4 cos x (1-cos^2(x)) cos x - 4 cos x + 4cos^3(x) -3cos x + 4cos^3(x) 4cos^3-3cosx, which is what the final we came up with before so just work it backwards into that ....

OpenStudy (umulas):

Ok, thanks!